OpenStudy (anonymous):

what is the slope of the line described by the equation -4x + 6y =12

OpenStudy (anonymous):

First we have to put the equation in the slope-intercept formula y = mx + b so -4x + 6y = 12 6y = 12 + 4x y = 2 + (2/3)x

OpenStudy (anonymous):

y = mx + b, so the coefficient in front of x is 2/3, therefore the slope of the line is 2/3

OpenStudy (anonymous):

good job brother. thanx.

OpenStudy (anonymous):

yw